# SuperCoder: Assembly Program Superoptimization with Large Language Models

### v2

@@ Abstract
Superoptimization is the task of transforming a program into a faster one while preserving its input–output behavior. In this work, we investigate whether large language models (LLMs) can serve as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers. We construct the first large-scale benchmark for this problem, consisting of 8,072 assembly programs averaging 130 lines, in contrast to prior datasets restricted to 2–15 straight-line, loop-free programs. We evaluate 23 LLMs on this benchmark and find that the strongest baseline, Claude-opus-4, achieves a 51.5% test-passing rate and a 1.43× average speedup over gcc -O3. To further enhance performance, we fine-tune models with reinforcement learning, optimizing a reward function that integrates correctness and performance speedup. Starting from Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup), the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup. Our results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly programs, establishing a foundation for future research in program performance optimization beyond compiler heuristics.
@@ 1 Introduction
Decades of research have tackled the problem of code optimization, giving rise to two main approaches. The first develops better algorithms for rule-based transformations in compilers (Wolf & Lam, 1991). However, given the vast space of possible transformations, compiler-optimized code is not guaranteed to be optimal and often leaves performance untapped (Center et al., 1971; Theodoridis et al., 2022). The second, superoptimization, develops search algorithms that directly explore the space of all possible programs, aiming to discover the correct variant with the best performance rather than relying on a fixed set of transformation rules (Schkufza et al., 2013).
Superoptimization is an aggressive form of program optimization that can outperform compiler-optimized code, yet existing literature has focused on very short, straight-line assembly programs without loops. Prior work has primarily relied on CPU-based search heuristics, which fail to scale to larger programs (Schkufza et al., 2013; Phothilimthana et al., 2016; Koenig et al., 2021b); available datasets include at most 15 lines of straight-line assembly (Koenig et al., 2021a).
@@
Unlike high-level programming languages (e.g., Python or C), large-scale, high-quality assembly datasets are scarce. As the first study in this direction, we construct a dataset of 8,072 assembly programs. Each instance includes input–output test cases and baseline assembly generated by the compiler at its highest optimization level (gcc -O3), which serves as the starting point for further optimization. In contrast, the datasets commonly used in the superoptimization community (Warren, 2013; Schkufza et al., 2013; Koenig et al., 2021a) are either extremely limited in size, containing only 25 programs, or consist of toy examples with 2 to 15 instructions, 74% of which have between 3 and 6 instructions and contain no loops. Our dataset is substantially larger, with assembly programs averaging 130 lines and including loops. Our test suites achieve 96.2% line and 87.3% branch coverage, demonstrating strong test quality. Our dataset represents a substantial step forward in scale for evaluating superoptimization techniques.
Beyond evaluating existing models, we also apply reinforcement learning (RL) for fine-tuning to further enhance their capabilities. We use widely adopted algorithms, including Proximal Policy Optimization (PPO) and Group Relative Policy Optimization (GRPO), to train an LLM with a reward function that integrates both correctness and performance speedup. Prior work on LLM-based performance optimization has explored alternative methodologies such as supervised fine-tuning (Shypula et al., 2023), chain-of-thought prompting (Liu et al., 2024c), agent-based frameworks (Wei et al., 2025b; c), and preference learning (Du et al., 2024). Our approach optimizes speedup explicitly in the reward function, making reinforcement learning well suited to superoptimization. To our knowledge, this is the first application of reward-based RL to LLMs for code performance optimization, with correctness and speedup jointly encoded in the objective.
We evaluate 23 LLMs on this task and find that the best-performing model, Claude-opus-4, achieves a 51.5% test-passing rate and an average speedup of 1.43× over the compiler-optimized baseline (gcc -O3). Our reinforcement learning approach is highly effective: starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves 61.4% correctness and a modest 1.10× speedup, the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup. We further analyze the code transformation patterns learned by the model.
@@
  • We evaluate 23 LLMs on the benchmark and show that RL-based training substantially improves performance: fine-tuning Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup) results in SuperCoder with 95.0% correctness and 1.46× speedup.
@@ 2 Related Work
Benchmarks for evaluating large language models (LLMs) on code generation from natural language specifications have received increasing attention. Notable examples include HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021), and more recent efforts (Liu et al., 2023b; Li et al., 2024; Xia et al., 2024b; Zhuo et al., 2024). In parallel, many models have been developed to enhance code generation capabilities, such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2022), InCoder (Fried et al., 2022), StarCoder (Li et al., 2023), DeepSeek-Coder (Guo et al., 2024), Code Llama (Roziere et al., 2023), and others (Hui et al., 2024; Wei et al., 2025e). Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair (Xia & Zhang, 2022; Xia et al., 2023b), software testing (Xia et al., 2023a; Deng et al., 2024), bug localization (Yang et al., 2024a), transpilation (Yang et al., 2024c; Bhatia et al., 2024), equivalence checking (Wei et al., 2025a), and synthesis (Wei et al., 2025d). SWE-bench (Jimenez et al., 2023) integrates these tasks into a benchmark for resolving real GitHub issues. Building on this, SWE-agent (Yang et al., 2024b) and subsequent works (Xia et al., 2024a; Wei et al., 2025f) employ an agent-based framework that leverages LLMs to improve the issue resolution process.
Recent work has also explored LLMs for improving program performance. CodeRosetta (TehraniJamsaz et al., 2024) targets automatic parallelization, such as translating C++ to CUDA. Other efforts focus on optimizing Python code for efficiency (Du et al., 2024; Liu et al., 2024c) or enabling self-adaptation (Huang et al., 2024), and improving C++ performance (Shypula et al., 2023). Of particular relevance are approaches to low-level code optimization (Wei et al., 2024; Ouyang et al., 2025). The LLM Compiler foundation models (Cummins et al., 2024; 2025) are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance. LLM-Vectorizer (Taneja et al., 2025) offers a formally verified solution for auto-vectorization, a specific compiler pass. In contrast, our work does not restrict the optimization type and uses test-case validation.
@@
More recently, LLMs have been explored as code optimizers (Shypula et al., 2023; Grubisic et al., 2024; Wei et al., 2024; 2025b; 2025c), with growing interest in reinforcement learning that guides generation through reward signals (Dou et al., 2024; Wei et al., 2025f). Rewards are often derived from unit-test correctness (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023a) or binary preference signals (Liu et al., 2024b; Du et al., 2024). To our knowledge, this is the first work to apply reinforcement learning to optimize code performance with LLMs, with concurrent efforts exploring CUDA kernel optimization (Li et al., 2025; Baronio et al., 2025).
@@ 3.1 Task Definition
The objective is to generate a candidate program $P^{\prime}$ that maximizes $\text{Speedup}(P^{\prime})$. Only programs in $\mathcal{S}(P)$ are eligible for speedup; any candidate that fails to compile into a binary or produces incorrect outputs is assigned a default speedup of 1. This reflects a practical fallback: when the generated program is invalid, the system can revert to the baseline $P$, compiled with gcc -O3, which defines the 1× reference performance. Although $\mathcal{S}(P)$ captures the correctness criteria, we do not restrict the LLM to generate only valid programs. Instead, the model produces arbitrary assembly code, and correctness is validated post hoc via compilation and test execution. We train an LLM using reinforcement learning (see Section˜3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
@@ 3.2 Dataset Construction
We construct our dataset using C programs from CodeNet (Puri et al., 2021), a large-scale corpus of competitive programming submissions. Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=gcc(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set. Since not all CodeNet problems include test inputs, we adopt those provided by prior work (Li et al., 2022) to define $x_{i}$, but discard their output labels. Instead, we regenerate each output $y_{i}$ by executing the original submission on input $x_{i}$, as many CodeNet programs are not accepted solutions, and even accepted ones do not reliably pass all test cases.
Given the scale of CodeNet, which contains over 8 million C and C++ submissions, we sample a subset for this study. To focus on performance-critical cases, we sample programs that exhibit the highest relative speedup from gcc -O0 (no optimization) to gcc -O3 (maximum optimization). Such a strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations. If an LLM can generate code that further improves upon gcc -O3, it suggests that the model can outperform the compiler’s “expert” solution. The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section˜4.
@@ 3.3 Reinforcement Learning
$t(P~),\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],\quad\operatorname{speedup}(s,a)$$t(\tilde{P})$,
@@ 4 Experimental Setup
We describe our dataset construction approach in Section˜3.2. Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation. The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in Table˜1, and additional analysis below.
Test Coverage.
@@ C Assembly
For each instance, we construct a prompt that includes the original C program along with the generated assembly using gcc -O3. Test cases are withheld from the model. The model is instructed to generate the optimized assembly code. We show the prompt template in Appendix A.3.
@@
We evaluate each model using both correctness and performance metrics. Compile pass is the percentage of problems for which the generated assembly compiles to binary executable successfully, while test pass is the percentage of problems where the compiled code passes all test cases. For a given problem, any single failed test case is considered a failure for the test pass metric. Both metrics are computed across the entire validation set. For performance, we measure the relative speedup over the gcc -O3 baseline. As defined in Section˜3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline. This reflects the practical setting where a system can fall back to the gcc -O3 output, resulting in no performance gain. We report the 25th, 50th (median), and 75th percentiles of speedup to capture distributional behavior, along with the average speedup over the entire evaluation set.
@@
We evaluate 23 state-of-the-art language models spanning a diverse range of architectures. Our benchmark includes frontier proprietary models such as gpt-4o (Achiam et al., 2023), o4-mini, gemini-2.0-flash-001 (Team et al., 2023), and claude-3.7-sonnet, as well as open-source families such as Llama (Touvron et al., 2023), DeepSeek (Liu et al., 2024a), and Qwen (Hui et al., 2024). In addition, we include models distilled from DeepSeek-R1 (Guo et al., 2025) based on Qwen and Llama. Finally, we evaluate recent compiler-oriented foundation models (Cummins et al., 2024; 2025), pretrained on assembly code and derived from Code Llama, with a design focus on compiler tasks (listed as llm-compiler in Table˜2). All open-source models are instruction-tuned.
@@
To ensure an accurate performance evaluation, we use hyperfine (hyp, 2025), a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions. For each program’s execution, we discard the first three runs and report the average runtime over the next ten runs.
@@
We implement our customized reinforcement learning reward functions within the verl framework (Sheng et al., 2024), which enables fine-tuning of LLMs using PPO and GRPO. As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section˜3.3. This environment provides the model with direct scalar feedback based on both functional correctness and execution performance.
@@
Among all evaluated models (see Table˜2), we select Qwen2.5-Coder-7B-Instruct for training due to its strong correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality. Training is performed on a single node with four A100 GPUs. Full hyperparameter settings are provided in Appendix A1.

### v3

@@ Abstract
Superoptimization is the task of transforming a program into a faster one while preserving its input–output behavior. In this work, we investigate whether large language models (LLMs) can serve as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers. We construct the first large-scale benchmark for this problem, consisting of 8,072 assembly programs averaging 130 lines, in contrast to prior datasets restricted to 2–15 straight-line, loop-free programs. We evaluate 23 LLMs on this benchmark and find that the strongest baseline, Claude-opus-4, achieves a 51.5% test-passing rate and a 1.43× average speedup over gcc -O3. To further enhance performance, we fine-tune models with reinforcement learning, optimizing a reward function that integrates correctness and performance speedup. Starting from Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup), the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup, with additional improvement enabled by Best-of-N sampling and iterative refinement. Our results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly programs, establishing a foundation for future research in program performance optimization beyond compiler heuristics.
Machine Learning, ICML
@@ 1 Introduction
Decades of research have tackled the problem of code optimization, giving rise to two main approaches. The first develops better algorithms for rule-based transformations in compilers (Wolf and Lam, 1991). However, given the vast space of possible transformations, compiler-optimized code is not guaranteed to be optimal and often leaves performance untapped (Center et al., 1971; Theodoridis et al., 2022). The second, superoptimization, develops search algorithms that directly explore the space of all possible programs, aiming to discover the correct variant with the best performance rather than relying on a fixed set of transformation rules (Schkufza et al., 2013).
Superoptimization is an aggressive form of program optimization that can outperform compiler-optimized code, yet existing literature has focused on very short, straight-line assembly programs without loops. Prior work has primarily relied on CPU-based search heuristics, which fail to scale to larger programs (Schkufza et al., 2013; Phothilimthana et al., 2016; Koenig et al., 2021a); available datasets include at most 15 lines of straight-line assembly (Koenig et al., 2021b).
@@
Unlike high-level programming languages (e.g., Python or C), large-scale, high-quality assembly datasets are scarce. As the first study in this direction, we construct a dataset of 8,072 assembly programs. Each instance includes input–output test cases and baseline assembly generated by the compiler at its highest optimization level (gcc -O3), which serves as the starting point for further optimization. In contrast, the datasets commonly used in the superoptimization community (Warren, 2013; Schkufza et al., 2013; Koenig et al., 2021b) are either extremely limited in size, containing only 25 programs, or consist of toy examples with 2 to 15 instructions without loops. Our dataset is substantially larger, with assembly programs averaging 130 lines and including loops. Our test suites achieve 96.2% line and 87.3% branch coverage, demonstrating strong test quality. Our dataset represents a substantial step forward in scale for evaluating superoptimization techniques.
Beyond evaluating existing models, we also apply reinforcement learning (RL) for fine-tuning to further enhance their capabilities. We use widely adopted algorithms, including Proximal Policy Optimization (PPO) and Group Relative Policy Optimization (GRPO), to train an LLM with a reward function that integrates both correctness and performance speedup. Prior work on LLM-based performance optimization has explored alternative methodologies such as supervised fine-tuning (Shypula et al., 2024), chain-of-thought prompting (Liu et al., 2024c), agent-based frameworks (Wei et al., 2025b, c), and preference learning (Du et al., 2024). Our approach optimizes speedup explicitly in the reward function, making reinforcement learning well-suited to superoptimization. To our knowledge, this is among the first applications of reward-based RL to LLMs for code performance optimization, with correctness and speedup jointly encoded in the objective.
We evaluate 23 LLMs on this task and find that the best-performing model, Claude-opus-4, achieves a 51.5% test-passing rate and an average speedup of 1.43× over the compiler-optimized baseline (gcc -O3). Our reinforcement learning approach is highly effective: starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves 61.4% correctness and a modest 1.10× speedup, the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup, with further improvement enabled by Best-of-N sampling and iterative refinement.
@@
  • We evaluate 23 LLMs on the benchmark and show that RL-based training substantially improves performance: fine-tuning Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup) results in SuperCoder with 95.0% correctness and 1.46× speedup, with further gains enabled by Best-of-N sampling and iterative refinement.
@@ 2 Related Work
Benchmarks for evaluating large language models (LLMs) on code generation from natural language specifications have received increasing attention. Notable examples include HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021), and more recent efforts (Liu et al., 2023b; Li et al., 2024; Xia et al., 2024b; Zhuo et al., 2024). In parallel, many models have been developed to enhance code generation capabilities, such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2022), InCoder (Fried et al., 2022), StarCoder (Li et al., 2023), DeepSeek-Coder (Guo et al., 2024), Code Llama (Roziere et al., 2023), and others (Hui et al., 2024; Wei et al., 2025e). Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair (Xia and Zhang, 2022; Xia et al., 2023b), software testing (Xia et al., 2023a; Deng et al., 2024), bug localization (Yang et al., 2024a), transpilation (Yang et al., 2024c; Bhatia et al., 2024), equivalence checking (Wei et al., 2025a), and synthesis (Wei et al., 2025d). SWE-bench (Jimenez et al., 2023) integrates these tasks into a benchmark for resolving real GitHub issues. Building on this, SWE-agent (Yang et al., 2024b) and subsequent works (Xia et al., 2024a; Wei et al., 2025f) employ an agent-based framework that leverages LLMs to improve the issue resolution process.
Recent work has also explored LLMs for improving program performance. CodeRosetta (TehraniJamsaz et al., 2024) targets automatic parallelization, such as translating C++ to CUDA. Other efforts focus on optimizing Python code for efficiency (Du et al., 2024; Liu et al., 2024c) or enabling self-adaptation (Huang et al., 2024), and improving C++ performance (Shypula et al., 2024). Of particular relevance are approaches to low-level code optimization (Wei et al., 2024; Ouyang et al., 2025). The LLM Compiler foundation models (Cummins et al., 2024, 2025) are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance. LLM-Vectorizer (Taneja et al., 2025) offers a formally verified solution for auto-vectorization, a specific compiler pass. In contrast, our work does not restrict the optimization type and uses test-case validation.
@@
More recently, LLMs have been explored as code optimizers (Shypula et al., 2024; Grubisic et al., 2024; Wei et al., 2024, 2025b, 2025c), with growing interest in reinforcement learning that guides generation through reward signals (Dou et al., 2024; Wei et al., 2025f). Rewards are often derived from unit-test correctness (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023a) or binary preference signals (Liu et al., 2024b; Du et al., 2024). To our knowledge, this is among the first works to apply reinforcement learning to optimize code performance with LLMs, with concurrent efforts exploring CUDA kernel optimization (Li et al., 2025; Baronio et al., 2025).
@@ 3.1 Task Definition
The objective is to generate a candidate program $P^{\prime}$ that maximizes $\text{Speedup}(P^{\prime})$. Only programs in $\mathcal{S}(P)$ are eligible for speedup; any candidate that fails to compile into a binary or produces incorrect outputs is assigned a default speedup of 1. This reflects a practical fallback: when the generated program is invalid, the system can revert to the baseline $P$, compiled with gcc -O3, which defines the 1× reference performance. Although $\mathcal{S}(P)$ captures the correctness criteria, we do not restrict the LLM to generate only valid programs. Instead, the model produces arbitrary assembly code, and correctness is validated post hoc via compilation and test execution. We train an LLM using reinforcement learning (see Section 3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
@@ 3.2 Dataset Construction
We construct our dataset using C programs from CodeNet (Puri et al., 2021), a large-scale corpus of competitive programming submissions. CodeNet is a well-established and widely used benchmark in the AI-for-code community (Li et al., 2022; Shypula et al., 2024). Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=gcc(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set. Since not all CodeNet problems include test inputs, we adopt those provided by prior work (Li et al., 2022) to define $x_{i}$, but discard their output labels. Instead, we regenerate each output $y_{i}$ by executing the original submission on input $x_{i}$, as many CodeNet programs are not accepted solutions, and even accepted ones do not reliably pass all test cases.
Given the scale of CodeNet, which contains over 8 million C and C++ submissions, we sample a subset for this study. To focus on performance-critical cases, we sample programs that exhibit the highest relative speedup from gcc -O0 (no optimization) to gcc -O3 (maximum optimization). Such a strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations. The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section 4. We also evaluate our method on an unbiased, randomly sampled dataset distribution in Appendix A.7.
@@ 3.3 Reinforcement Learning
$\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],$
$t(P~),\operatorname{speedup}(s,a)$$t(\tilde{P})$,
@@
3.4 Best-of-N Sampling, Supervised Fine-Tuning, and Iterative Refinement
Best-of-N Sampling.
Generating multiple candidate programs and selecting the strongest one is a well-established strategy for improving code generation quality (Li et al., 2022; Ehrlich et al., 2025). In our setting, the best candidate refers to the program that is correct while achieving the fastest execution time. Best-of-N sampling is an inference-time technique that can boost performance, but it incurs additional cost because each candidate is tested at runtime.
Supervised Fine-Tuning.
To obtain training targets for supervised fine-tuning, we require reference solutions to the superoptimization task. However, superoptimization is inherently open-ended: beyond the compiler baseline, there is no unique ground-truth program, and multiple distinct solutions may exist. We apply best-of-8 sampling with the base model and treat the highest-quality candidate for each instance as the ground truth. We then fine-tune the model using LoRA (Hu et al., 2022).
Iterative Refinement.
Iterative refinement is a complementary inference-time technique that can be applied to any model to further improve its outputs. After each trial, we feed back the model’s previous attempt: if the generated program fails to compile or fails any test cases, we include the corresponding compiler errors or test failures in the next prompt; if the model produces a correct program, we also include that successful attempt as part of the prompt.
@@ C Assembly
We describe our dataset construction approach in Section 3.2. Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation. The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in LABEL:tab:dataset, and additional analysis below.
Test Coverage.
@@
For each instance, we construct a prompt that includes the original C program along with the generated assembly using gcc -O3. Test cases are withheld from the model. We show the prompt template in Appendix A.3.
@@
We evaluate each model using both correctness and performance metrics. Compile pass is the percentage of problems for which the generated assembly compiles to binary executable successfully, while test pass is the percentage of problems where the compiled code passes all test cases. For a given problem, any single failed test case is considered a failure for the test pass metric. Both metrics are computed across the entire validation set. For performance, we measure the relative speedup over the gcc -O3 baseline. As defined in Section 3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline. This reflects the practical setting where a system can fall back to the gcc -O3 output, resulting in no performance gain. We report the 25th, 50th (median), and 75th percentiles of speedup to capture distributional behavior, along with the average speedup over the entire evaluation set.
@@
We evaluate 23 state-of-the-art language models spanning a diverse range of architectures. Our benchmark includes frontier proprietary models such as gpt-4o (Achiam et al., 2023), o4-mini, gemini-2.0-flash-001 (Team et al., 2023), and claude-3.7-sonnet, as well as open-source families such as Llama (Touvron et al., 2023), DeepSeek (Liu et al., 2024a), and Qwen (Hui et al., 2024). In addition, we include models distilled from DeepSeek-R1 (Guo et al., 2025) based on Qwen and Llama. Finally, we evaluate recent compiler-oriented foundation models (Cummins et al., 2024, 2025), pretrained on assembly code and derived from Code Llama, with a design focus on compiler tasks (listed as llm-compiler in Table 2). All open-source models are instruction-tuned.
@@
To ensure an accurate performance evaluation, we use hyperfine (27), a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions. For each program’s execution, we discard the first three runs and report the average runtime over the next ten runs.
@@
We implement our customized reinforcement learning reward functions within the verl framework (Sheng et al., 2024), which enables fine-tuning of LLMs using PPO and GRPO. As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section 3.3. This environment provides the model with direct scalar feedback based on both functional correctness and execution performance.
@@
Among all evaluated models (see Table 2), we select Qwen2.5-Coder-7B-Instruct for training due to its strong correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality. Training is performed on a single node with four A100 GPUs. Hyperparameter settings are in Appendix A.2.